OpenStudy (anonymous):

My answer is A Solve 6x = 24 (2 points) Answer x ≈ 30.84 x ≈ 1.77 x ≈ 0.56 x ≈ 0.23

jimthompson5910 (jim_thompson5910):

is it \[\Large 6^x = 24\]

OpenStudy (anonymous):

yeah @jim_thompson5910 I believe so. if it would be: 6x = 24 then x = 4

OpenStudy (anonymous):

Nope, it is x = 1.77

OpenStudy (anonymous):

it's 6^x

OpenStudy (anonymous):

Solving Exponential Equations with Unequal Bases

OpenStudy (anonymous):

It's not 4....

jimthompson5910 (jim_thompson5910):

\[\Large 6^x = 24\] \[\Large \ln(6^x) = \ln(24)\] \[\Large x*\ln(6) = \ln(24)\] \[\Large x = \frac{\ln(24)}{\ln(6)}\] \[\Large x \approx 1.77370561446909\] \[\Large x \approx 1.77\]

OpenStudy (anonymous):

was it seriously THAT easy????

OpenStudy (anonymous):

I'm bad at roots and such, but plug and chug works here. Plug in 1.77 to yield 23.8

OpenStudy (anonymous):

thank you so much guys